Please help me factor the polynomial 2y^4-13(y^3)z-132(y^2)(z^2)
Found 2 solutions by uma, stanbon:
Answer by uma(370)


2y^4 - 13y^3z - 132y^2z^2
= y^2 (2y^2 - 13yz -132z^2) [removing the common factor y^2 from all the terms]
= y^2(2y^2 - 24yz + 11yz - 132z^2) [splitting the middle term]
= y^2[2y(y - 12z) + 11z(y - 12z)]
= y^2[(y - 12z)(2y + 11z)]
= y^2(y - 12z)(2y+ 11z)

Good Luck!!!

Answer by stanbon(48569)

factor the polynomial
2y^4-13(y^3)z-132(y^2)(z^2)
=Y^2(2y^2-13yz-132z^2)
=y^2(2y+11z)(y-12z)
Cheers,
Stan H.